OpenStudy (anonymous):

simplify cube rt 6xy^3 (times) cube rt 18x^4y

OpenStudy (anonymous):

\[\sqrt[3]{6xy^3}*\sqrt[3]{18x^4y}=\sqrt[3]{108x^5y^4}=\sqrt[3]{27x^3y^3}*\sqrt[3]{4x^2y}=3xy*\sqrt[3]{4x^2y}\]